# Parallel computations for Metropolis Markov chains with Picard maps

### v1

@@ 7 Extensions
We outline here some extensions of this work, which are left for future research:
  • The Online Picard recursion (3) sets $X^{(j+1)}_{i}$ for $i\geq U^{(j)}$ to a constant value equal to the last coordinate of the Picard map $\Phi$. While this choice is simple and reasonable, it does not exploit the intrinsic dynamics of the underlying Markov chain and the noise variables beyond the time index $U^{(j)}$. More elaborate recursions are possible, combining Picard maps with inexpensive predictive functions that estimate the values of the Markov chain at those steps where the Picard map is not been applied yet. An example of such schemes is provided by the (stochastic) Parareal algorithm (Lions et al., 2001; Pentland et al., 2022).
  • By analyzing the family of approximate measures $\pi_{r}$ for $r>0$ in Section 5.1, it may be possible to establish the parallel round complexity of the Approximate Online Picard algorithm. Our numerical results suggest that this algorithm can achieve a parallel round complexity of order $\mathcal{O}(1)$ using $\mathcal{O}(d)$ parallel processors.
  • The proof of Proposition 1 shows that $L^{(j+1)}=L^{(j)}+G^{(j)}$ with $G^{(j)}=G(\bar{X}^{(j)},\bar{W}^{(j)})$ and $G$ a deterministic function. Assuming that $(\bar{X}^{(j)},\bar{W}^{(j)})_{j\in\mathbb{N}}$ has a unique invariant measure $\rho$, then we can interpret $\mathbb{E}_{\rho}(G(\bar{X},\bar{W}))$ as the average speedup relative to the sequential algorithm in stationarity and $N/\mathbb{E}_{\rho}(G(\bar{X},\bar{W}))$ is the expected number of parallel iterations of the Online Picard algorithm for simulating the Markov chain $X_{0:N}$. This suggest that trying to characterize explicitly the invariant distribution $\rho$ might be a promising way to improve the analysis of the Online Picard algorithm, e.g. going beyond classical log-concavity or obtaining more explicit expressions.
@@ Acknowledgments
Both authors acknowledge support from the European Union (ERC), through the Starting grant ‘PrSc-HDBayLe’, project number 101076564.
Appendix A Proofs for Random Walk Metropolis
A.1 Background lemmas
@@
Proof.
Without loss of generality, let $a\leq b$. For the cases $a,b\leq 0$ and $a\leq 0,b\geq 0$, Lemma 1 is trivial, while for $a,b\geq 0$, it follows from the fact that the density function of an exponential random variable is non-increasing. ∎
Lemma 2.
@@
Proof.
By rotational invariance of $\mathcal{N}(0,\sigma^{2}I_{d})$, we have $\mathbb{E}[|\langle v,Z\rangle|=\|v\|\mathbb{E}[|Z_{1}|]$, where $|Z_{1}|$ follows a one-dimensional folded normal distribution and has mean $\sqrt{\frac{2}{\pi}}\sigma$. ∎
Lemma 3.
@@
Proof.
$\mathbb{E}[Z^{\top}MZ]=\sum_{i,j}\mathbb{E}[Z_{i}Z_{j}]\mathbb{E}[M_{i,j}]=\sum_{i}\mathbb{E}[Z_{i}^{2}]\mathbb{E}[M_{i,i}]=\sigma^{2}\mathrm{Tr}(\mathbb{E}[M])$. ∎
Lemma 4.
@@
where we used Fubini-Tonelli’s Theorem in the third equality. ∎
A.2 Contraction of the Picard map: key lemmas
@@
Lemma 5.
Under Assumptions 2-3, for every $x,y\in\mathcal{X}$,
@@
Proof.
By (8), under Assumptions 2-3, we have that
@@
with $U\sim\text{Unif}([0,1])$ and $Z\sim\mathcal{N}(0,h^{2}/(Ld)I_{d})$. Since $E=-\log(U)\sim\mathcal{E}(1)$, by Lemma 1 we have
@@
For the first term, by $Z\sim\mathcal{N}(0,c^{2}/(Ld)I_{d})$, Lemma 2 and the $L$-smoothness of $V$, we have
$\mathbb{E}[|\Delta_{1}|]=\sqrt{\frac{2}{\pi}}\frac{hL^{1/2}}{d^{1/2}}\|\nabla V(y)-\nabla V(x)\|\leq\sqrt{\frac{2}{\pi}}\frac{h}{(Ld)^{1/2}}\|x-y\|^{2}.$
For the second term, by the strong Hessian-Lipschitz condition in Assumption 2 and Lemma 3, we have
@@
as desired. ∎
Lemma 6.
Under Assumption 3 and for all $x,y\in\mathcal{X}^{K+1}$ with $x_{0}=y_{0}$, $w_{0}\in\mathcal{W}$ and $1<i\leq d$,
@@
with $W_{1:K-1}\sim\nu^{\otimes(K-1)}$, $\nu^{\otimes p}$ being product measure of $\nu$ repeated $p$ times, and $\delta(t)$ as in Theorem 1. The inequality (16) holds also when $x$ and $y$ are replaced by random variables $X,Y$ on $\mathcal{X}^{K+1}$, which can depend on $W$.
The proof of Lemma 6 relies on the following bounds.
Lemma 7.
@@
Proof.
By Wainwright [2019, Theorem 6.1], for all $t\geq 0,\,i\leq d$,
@@
for its maximum eigenvalue. Combining this result and Lemma 4, we have
@@
with $\delta(d)=\frac{5}{3}\exp(-3d/2)$. ∎
Proof of Lemma 6.
@@
[…]
@@
with $\Delta B=(\Delta B_{1},\dots,\Delta B_{i-1})^{T}$ and $W\sim\hbox{Wishart}(I_{i-1},d)$. Here =$\stackrel{{\scriptstyle d}}{{=}}$ denotes equality in distribution. Combining (17)-(18) and using Lemma 7 we have
@@
with $\delta(d)=\frac{5}{3}\exp(-3d/2)$. The result follow by $\mathbb{E}[\|\Delta B\|^{2}]=\sum_{\ell=1}^{i-1}\mathbb{P}(B(x_{\ell},U_{\ell},Z_{\ell})\neq B(y_{\ell},U_{\ell},Z_{\ell}))$. ∎
@@
Lemma 8.
Under the assumptions of Theorem 1, we have
@@
with $c_{0}$, $\delta(d)$ as in Theorem 1.
Proof.
By Lemma 5 and Jensen’s inequality we have
@@
By Lemma 6 we have that
@@
which implies the desired result after noting that $\max_{\ell\leq i}$ is a non decreasing operator in $i$. ∎
A.3 Proof of Theorem 1
Lemma 9.
@@ A.3 Proof of Theorem 1
Proof.
Let $g(a)=b\sqrt{a+\epsilon}$. By monotonicity of $g$, we can assume without loss of generality that $a_{0}=1$ and $a_{j+1}=g(a_{j})$ for all $j\geq 1$. The function $g$ is concave and its unique fixed point on $[0,\infty)$ satisfying $a^{*}=g(a^{*})$ is
@@
If $a_{0}=1\leq a^{*}$ then $a_{j}\leq a^{*}\leq b^{2}+\epsilon\leq b^{2}+\epsilon+2^{-j}$ for all $j\geq 0$, as desired. Consider now $a_{0}=1\geq a^{*}$. Since $g^{\prime}$ is strictly decreasing on $[0,\infty)$, for all $a\geq a^{*}$ we have $20<g^{\prime}(a)\leq g^{\prime}(a^{*})\leq g^{\prime}(b^{2}-\epsilon)=1/$ 2. Thus, $g$ is a strict contraction on $[a^{*},\infty)$ with Lipschitz constant smaller than $1/2$ and, by Banach’s fixed point theorem, we have $|a_{j}-a^{*}|\leq 2^{-j}|a_{0}-a^{*}|$ for all $j\geq 0$. By $a_{0}=1\geq a^{*}$, we have $|a_{0}-a^{*}|\leq a_{0}=1$ and thus we deduce $a_{j}\leq a^{*}+2^{-j}\leq b^{2}+\epsilon+2^{-j}$ as desired. ∎
Proof of Theorem 1.
. Follows directly by Lemma 8 and Lemma 9 with $b=\sqrt{c_{0}i/d}$, $\epsilon=\delta(d)$, and $c_{0},\delta(d)$ as in Theorem 1. ∎
A.4 Proof of Corollary 1
@@
Applying the union bound and the bound in Theorem 1, for $K\leq b\sqrt{d},b>0$, it follows
@@
This is less than $\epsilon$ if $c_{0}b^{2}=\frac{\epsilon}{2}$ and $b\sqrt{d}(\delta(d)+2^{-j})\leq\epsilon/2$. The first condition is achieved by setting $b=\frac{\epsilon}{\sqrt{2c_{0}}}$. With this choice, the latter condition is always satisfied for $h\geq 1,\,d\geq 1$ since $(h,d)\mapsto\sqrt{\frac{d}{2c_{0}}}(\delta(d)+2^{-j})-\frac{1}{2}$ is negative on $[1,\infty)^{2}$.
A.5 Proof of Theorem 2
@@
Proposition 5.
Under Assumption 2-3, $h\geq 1$, for all $\epsilon>0$, $K\leq\sqrt{\epsilon^{2}d/(2c_{0})}$, $x_{0}\in\mathcal{X}$, any $j\in\mathbb{N}$ and $p=\lceil\log(d)\rceil$, we have
@@
The proof of Proposition 5 requires the following two lemmas. For any $i\in\mathbb{N}$, define $\tilde{G}_{i}:=\sum_{\ell=0}^{p-1}G^{(pi+\ell)}$, with $(G^{(i)})_{i=0,1,\dots}$ defined in (7).
Lemma 10.
@@
where $F_{X}(t)$ denotes the distribution function of $X$ evaluated at $t$ and the last inequality follows since $u\mapsto F_{P_{i+1}}(t-u)$ is decreasing. ∎
Proof of Proposition 5.
As shown in the proof of Proposition 1, we have $L^{(j+1)}=\sum_{i=0}^{j}G^{(i)}$. Let $\tilde{L}_{i}:=L^{(pi)}=\sum_{\ell=0}^{i-1}\tilde{G}_{\ell}$. Note that $\{\tilde{L}_{i}=l\}\in\sigma(X_{0},W_{0:l})$ where $\sigma(X)$ denotes the sigma-algebra generated by the random variable $X$.
@@
where $x_{i}=x_{0}+\sum_{\ell=1}^{i-1}f(x_{\ell},w_{\ell}),\,i\geq 1$ and the last inequality follows from Corollary 1 and from $\tilde{G}_{i}\geq p$ almost surely, which holds by construction. Combining the above with Lemmas 10-11 yields the desire result. ∎
Proof of Theorem 2.
Fix $\epsilon=1/2$. By Proposition 5, for all $j\in\mathbb{N}$, $p=\lceil\log(d)\rceil$, $K\leq\sqrt{d/(8c_{0})}$, we have
@@
with $(P_{i})_{i=1}^{j}$ as in Proposition 5, $\mu=\mathbb{E}[P_{1}]=(1-\epsilon)K=K/2$. By Hoeffding’s inequality, we have
@@
$\leq\exp\left(-\frac{2(jK/2-n)^{2}}{jK^{2}}\right)=\exp\left(-j\left(1-\frac{2n}{jK}\right)\right)$
for any $j>2n/K$. In particular, if $j\geq\max\{\frac{4n}{K},2\log(1/\delta)\}$, then $j\left(1-\frac{2n}{jK}\right)\geq 2\log(1/\delta)\left(1-\frac{1}{2}\right)=\log(1/\delta)$ and the above exponential bound is less than $\delta$. ∎
A.6 Proof of Corollary 2
Let $(X_{n})_{n=0,1,2,\dots}$ be the RWM trajectory, i.e. the exact solution of the recursion in (1). Under Assumptions 2-3, $mI_{d}\preccurlyeq\nabla^{2}V$, $h=1$ and $X_{0}\sim\mathcal{N}(x^{\star},L^{-1}I_{d})$, we can apply Theorem 49 and Remark 50 of Andrieu et al. [2024] to deduce $\sqrt{\chi^{2}(\mathcal{L}(X_{n_{mix}}),\pi)}\leq\epsilon$ with $n_{mix}=\mathcal{O}(\frac{L}{m}d)$, where $n_{mix}$ depends on $\epsilon$, $d$, $L/m$ and is explicitly provided in Andrieu et al. [2024, Rmk.50]. Consider running the Online Picard algorithm with $K=\sqrt{d/(8c_{0})}$ for
@@ A.6 Proof of Corollary 2
parallel iterations and returning $Y=X^{(J)}_{\min\{n_{mix},L^{(J)}\}}$ as output. When $n_{mix}\leq L^{(J)}$, we have $Y=X^{(J)}_{n_{mix}}=X_{n_{mix}}$. By Theorem 2, $\mathbb{P}(Y=X_{n_{mix}})\geq\mathbb{P}(L^{(J)}\geq n_{mix})\geq 1-\epsilon/2$. Therefore
@@ A.7 Proof of Proposition 2
Lemma 12.
Under Assumption 2 with $mI_{d}\preccurlyeq\nabla^{2}V$, we have
$\mathbb{E}[|A_{n}-B_{n}|]$ $\leq 2L(C+d),$ for all $\hbox{for all }n=1,2,\dots\,,$ (19)
$\limsup_{n}\mathbb{P}(|A_{n}|\leq c)$ $=0$ for $\hbox{for all }n=1,2,\dots\,,\,c\in(0,\infty).$ (20)
Proof.
Let $A=V(x+z)-V(x)$, $B=V(y+z)-V(y)$, for some $x,y,z\in\mathcal{X}$. By $L$-smoothness of $V$, we have that, for some $t_{1},t_{2}\in(0,1)$ depending on $x,y,z$,
@@
By rotational invariance $Z$, the right hand-side of (21) depends on $\nabla V(x_{n})$ only through its norm. By $m$-convexity of $V$, $\|\nabla V(x_{n})\|\to\infty$ as $|x_{n}|\to\infty$ and, consequently, the right hand-side of (21) goes to 0, proving the desired result in (20). ∎
Lemma 13.
Under Assumption 2 with $mI_{d}\preccurlyeq\nabla^{2}V$, we have
@@
By Lemma 12, $\limsup_{n\to\infty}\mathbb{P}(A_{n}>c\cap B_{n}>c)=0$ and
@@
which goes to 0 as $n\to\infty$ by Lemma 13. ∎
A.8 Proof of Proposition 4
By Markov inequality and the bound in Theorem 1, we have for $K\leq\lfloor bd\rfloor$, $b=\epsilon r/(3c_{0})$,

### v2

@@ 7 Extensions
We outline here some extensions of this work, which are left for future research. First, the Online Picard recursion (3) sets $X^{(j+1)}_{i}$ for $i\geq U^{(j)}$ to a constant value equal to the last coordinate of the Picard map $\Phi$. While this choice is simple and reasonable, it does not exploit the intrinsic dynamics of the underlying Markov chain and the noise variables beyond the time index $U^{(j)}$. More elaborate recursions are possible, combining Picard maps with inexpensive predictive functions that estimate the values of the Markov chain at those steps where the Picard map is not been applied yet. An example of such schemes is provided by the (stochastic) Parareal algorithm (Lions et al., 2001; Pentland et al., 2022). It would be interesting to explore and characterize the gains obtained by these approaches in our high-dimensional zeroth-order context. Second, by analyzing the family of approximate measures $\pi_{r}$ for $r>0$ in Section 5, it may be possible to establish the parallel round complexity of the Approximate Online Picard algorithm. Our numerical results suggest that this algorithm can achieve a parallel round complexity of order $\mathcal{O}(1)$ using $\mathcal{O}(d)$ parallel processors. Finally, Proposition D.1 in the supplementary material suggests that trying to characterize explicitly the invariant distribution of the Markov chain $(X^{(j)}_{L^{(j)}:L^{(j)}+K},W_{L^{(j)}:L^{(j)}+K-1})_{j\in\mathbb{N}}$ might be a promising way to improve the analysis of the Online Picard algorithm, e.g. going beyond classical log-concavity or obtaining more explicit expressions, or quantifying the error made by the approximate Picard version.
@@ Acknowledgments
Both authors acknowledge support from the European Research Council (ERC), through the Starting grant ‘PrSc-HDBayLe’, project number 101076564. Both authors are grateful to Florian Maire for guiding us through the black-box code used for the application in Section 6.8.
Supplementary Materials for “Parallel computations for Metropolis Markov chains with Picard maps” by S. Grazzi and G. Zanella
Appendix A Assumption 3.1 for Bayesian Logistic Regression
Here we verify Assumption 3.1 for the logistic regression with Gaussian prior $\mathcal{N}_{d}(0,\Lambda^{-1}_{0})$. Assume for $i=1,2,\dots,n$, $Y_{i}\mid A_{i}\overset{\text{i.i.d.}}{\sim}\mbox{Ber}(\psi(\langle A_{i},x\rangle)),$ with $\psi(u)=(1+e^{-u})^{-1}$ and unknown parameter $x\in\mathbb{R}^{d}$, response variables $Y_{i}\in\{0,1\}$ and covariates $A_{i}\in\mathbb{R}^{d}$. Then, the negative log-posterior takes the form
$V(x)=\frac{1}{2}x^{\top}\Lambda_{0}x+\sum_{i=0}^{n}\left(y_{i}\log(\psi(A_{i}^{\top}x))+(1-y_{i})\log(1-\psi(A_{i}^{\top}x))\right)$
with Hessian
$\nabla^{2}V(x)=\Lambda_{0}+\sum_{i=1}^{n}\psi(A_{i}^{\top}x)(1-\psi(A_{i}^{\top}x))A_{i}A_{i}^{\top}.$
Since $\psi(A_{i}^{\top}x)(1-\psi(A_{i}^{\top}x))\leq 1/4$, we have that
$\nabla^{2}V(x)\preccurlyeq\Lambda_{0}+\frac{1}{4}\sum_{i=1}^{n}A_{i}A_{i}^{\top},$
thus $V$ is $L$-smooth with
$L:=\lambda_{max}(\Lambda_{0}+\frac{1}{4}\sum_{i=1}^{n}A_{i}A_{i}^{\top}).$
Define $w(t):=\psi(t)(1-\psi(t))$. One can check that $w$ is Lipschitz with constant $C=1/(6\sqrt{3})$. Since $\nabla^{2}V(y)-\nabla^{2}V(x)=\sum_{i=1}^{n}(w(A_{i}^{\top}y)-w(A_{i}^{\top}x))A_{i}A_{i}^{\top}$, we have
$\|y-x\|M\preccurlyeq\nabla^{2}V(y)-\nabla^{2}V(x)\preccurlyeq\|y-x\|M$
with $M=C\sum_{i=1}^{n}\|A_{i}\|A_{i}A_{i}^{\top}$. Thus, Assumption 3.1 holds with $\gamma=\mathrm{Tr}(M)/(L^{3/2}d^{1/2}),$ and $\mathrm{Tr}(M)=C\sum_{i=1}^{n}\|A_{i}\|^{3}$. Note that, if one further assumes that $A_{i}\sim\mathcal{N}(0,I_{d}/d)$ (as in Section 6.4), $\Lambda_{0}=\alpha I_{d}$ and $n/d\to\rho<\infty$, then, by standard high-probability bounds for Gaussians, one can easily show that $\mathrm{Tr}(M)=nC(1+o_{\mathbb{P}}(1))$, $L=\alpha+\frac{1}{4}(1+\sqrt{\rho})^{2}+o_{\mathbb{P}}(1)$, where $X_{d}=o_{\mathbb{P}}(a_{d})$ means $X_{d}/a_{d}\to 0$ in probability. Thus
$\gamma=C\frac{n}{L^{3/2}d^{1/2}}(1+o_{\mathbb{P}}(1))\,.$
Appendix B Proofs for Random Walk Metropolis
B.1 Background lemmas
@@
Proof B.2.
Without loss of generality, let $a\leq b$. For the cases $a,b\leq 0$ and $a\leq 0,b\geq 0$, Lemma B.1 is trivial, while for $a,b\geq 0$, it follows from the fact that the density function of an exponential random variable is non-increasing.
Lemma B.3.
@@
Proof B.4.
By rotational invariance of $\mathcal{N}(0,\sigma^{2}I_{d})$, we have $\mathbb{E}[|\langle v,Z\rangle|=\|v\|\mathbb{E}[|Z_{1}|]$, where $|Z_{1}|$ follows a one-dimensional folded normal distribution and has mean $\sqrt{\frac{2}{\pi}}\sigma$.
Lemma B.5.
@@
Proof B.6.
$\mathbb{E}[Z^{\top}MZ]=\sum_{i,j}\mathbb{E}[Z_{i}Z_{j}]\mathbb{E}[M_{i,j}]=\sum_{i}\mathbb{E}[Z_{i}^{2}]\mathbb{E}[M_{i,i}]=\sigma^{2}\mathrm{Tr}(\mathbb{E}[M])$.
Lemma B.7.
@@
where we used Fubini-Tonelli’s Theorem in the third equality.
B.2 Contraction of the Picard map: key lemmas
@@
Lemma B.9.
Under Assumptions 3.1-3.2, for every $x,y\in\mathcal{X}$,
@@
Proof B.10.
By (5), under Assumptions 3.1-3.2, we have that
@@
with $U\sim\text{Unif}([0,1])$ and $Z\sim\mathcal{N}(0,h^{2}/(Ld)I_{d})$. Since $E=-\log(U)\sim\mathcal{E}(1)$, by Lemma B.1 we have
@@
For the first term, by $Z\sim\mathcal{N}(0,h^{2}/(Ld)I_{d})$, Lemma B.3 and the $L$-smoothness of $V$, we have
$\mathbb{E}[|\Delta_{1}|]=\sqrt{\frac{2}{\pi}}\frac{h}{\sqrt{Ld}}\|\nabla V(y)-\nabla V(x)\|\leq\sqrt{\frac{2}{\pi}}h\sqrt{\frac{L}{d}}\|x-y\|^{2}.$
For the second term, by the strong Hessian-Lipschitz condition in Assumption 3.1 and Lemma B.5, we have
@@
as desired.
Lemma B.11.
Under Assumption 3.2 and for all $x,y\in\mathcal{X}^{K+1}$ with $x_{0}=y_{0}$, $w_{0}\in\mathcal{W}$ and $1<i\leq d$,
@@
with $W_{1:K-1}\sim\nu^{\otimes(K-1)}$, $\nu^{\otimes p}$ being product measure of $\nu$ repeated $p$ times, and $\delta(t)$ as in Theorem 3.1. The inequality (14) holds also when $x$ and $y$ are replaced by random variables $X,Y$ on $\mathcal{X}^{K+1}$, which can depend on $W$.
The proof of Lemma B.11 relies on the following bounds.
Lemma B.12.
@@
Proof B.13.
By Wainwright (2019, Theorem 6.1), for all $t\geq 0,\,i\leq d$,
@@
for its maximum eigenvalue. Combining this result and Lemma B.7, we have
@@
with $\delta(d)=\frac{5}{3}\exp(-3d/2)$.
Proof B.14 (of Lemma B.11).
@@
where $\Delta B_{\ell}=B(x_{\ell},U_{\ell},Z_{\ell})-B(y_{\ell},U_{\ell},Z_{\ell})$, and (15) is independent of $W_{0}$. Here $U_{1},U_{2},\dots,U_{i-1}\overset{\mathrm{i.i.d}}{\sim}\text{Unif}([0,1])$ and $Z_{1},Z_{2},\dots,Z_{i-1}\overset{\text{i.i.d}}{\sim}\mathcal{N}(0,h^{2}/(Ld))$. Also,
@@
with $\Delta B=(\Delta B_{1},\dots,\Delta B_{i-1})^{T}$ and $W\sim\hbox{Wishart}(I_{i-1},d)$. Here $\stackrel{{\scriptstyle d}}{{=}}$ denotes equality in distribution. Combining (15)-(16) and using Lemma B.12 we have
@@
with $\delta(d)=\frac{5}{3}\exp(-3d/2)$. The result follow by $\mathbb{E}[\|\Delta B\|^{2}]=\sum_{\ell=1}^{i-1}\mathbb{P}(B(x_{\ell},U_{\ell},Z_{\ell})\neq B(y_{\ell},U_{\ell},Z_{\ell}))$.
@@
Lemma B.15.
Under the assumptions of Theorem 3.1, we have
@@
with $c_{0}$, $\delta(d)$ as in Theorem 3.1.
Proof B.16.
By Lemma B.9 and Jensen’s inequality we have
@@
By Lemma B.11 we have that
@@
which implies the desired result after noting that $\max_{\ell\leq i}$ is a non decreasing operator in ii.
B.3 Proof of Theorem 3.1
Lemma B.17.
@@
Proof B.18.
Let $g(a)=b\sqrt{a+\epsilon}$. By monotonicity of $g$, we can assume without loss of generality that $a_{0}=1$ and $a_{j+1}=g(a_{j})$ for all $j\geq 1$. The function $g$ is concave and its unique fixed point on $0,\infty)$ satisfying $a^{*}=g(a^{*})$ is
@@
If $a_{0}=1\leq a^{*}$ then $a_{j}\leq a^{*}\leq b^{2}+\epsilon\leq b^{2}+\epsilon+2^{-j}$ for all $j\geq 0$, as desired. Consider now $a_{0}=1\geq a^{*}$. Since $g^{\prime}$ is strictly decreasing on $0,\infty)$, for all $a\geq a^{*}$ we have $0<g^{\prime}(a)\leq g^{\prime}(a^{*})\leq g^{\prime}(b^{2}-\epsilon)=1/2$. Thus, $g$ is a strict contraction on $a^{*},\infty)$ with Lipschitz constant smaller than $1/2$ and, by Banach’s fixed point theorem, we have $|a_{j}-a^{*}|\leq 2^{-j}|a_{0}-a^{*}|$ for all $j\geq 0$. By $a_{0}=1\geq a^{*}$, we have $|a_{0}-a^{*}|\leq a_{0}=1$ and thus we deduce $a_{j}\leq a^{*}+2^{-j}\leq b^{2}+\epsilon+2^{-j}$ as desired.
Proof B.19 (of Theorem 3.1).
. Follows directly by Lemma B.15 and Lemma B.17 with $b=\sqrt{c_{0}i/d}$, $\epsilon=\delta(d)$, and $c_{0},\delta(d)$ as in Theorem 3.1.
B.4 Proof of Corollary 3.4
@@
Applying the union bound and the bound in Theorem 3.1, for $K\leq b\sqrt{d},b>0$, it follows
@@
This is less than $\epsilon$ if $c_{0}b^{2}=\frac{\epsilon}{2}$ and $b\sqrt{d}(\delta(d)+2^{-j})\leq\epsilon/2$. The first condition is achieved by setting $b=\frac{\epsilon}{\sqrt{2c_{0}}}$. With this choice, the latter condition is always satisfied for $h\geq 1,\,d\geq 1$ since $(h,d)\mapsto\sqrt{\frac{d}{2c_{0}}}(\delta(d)+2^{-j})-\frac{1}{2}$ is negative on $1,\infty)^{2}$.
B.5 Proof of Theorem 3.5
@@
Proposition B.20.
Under Assumption 3.1-3.2, $h\geq 1$, for all $\epsilon>0$, $K\leq\sqrt{\epsilon^{2}d/(2c_{0})}$, $x_{0}\in\mathcal{X}$, any $j\in\mathbb{N}$ and $p=\lceil\log(d)\rceil$, we have
@@
The proof of Proposition B.20 requires the following two lemmas. For any $i\in\mathbb{N}$, define $\tilde{G}_{i}:=\sum_{\ell=0}^{p-1}G^{(pi+\ell)}$, with $(G^{(i)})_{i=0,1,\dots}$ defined in (28) below, $\delta_{c}(\cdot)$ being the Dirac measure centered at $c$.
Lemma B.21.
@@
where $F_{X}(t)$ denotes the distribution function of $X$ evaluated at $t$ and the last inequality follows since $u\mapsto F_{P_{i+1}}(t-u)$ is decreasing.
Proof B.25 (of Proposition B.20).
As also shown in the proof of Proposition D.1 below, we have $L^{(j+1)}=\sum_{i=0}^{j}G^{(i)}$. Let $\tilde{L}_{i}:=L^{(pi)}=\sum_{\ell=0}^{i-1}\tilde{G}_{\ell}$. Note that $\{\tilde{L}_{i}=l\}\in\sigma(X_{0},W_{0:l})$ where $\sigma(X)$ denotes the sigma-algebra generated by the random variable $X$.
@@
where $x_{i}=x_{0}+\sum_{\ell=1}^{i-1}f(x_{\ell},w_{\ell}),\,i\geq 1$ and the last inequality follows from Corollary 3.4 and from $\tilde{G}_{i}\geq p$ almost surely, which holds by construction. Combining the above with Lemmas B.21-B.23 yields the desire result.
Proof B.26 (of Theorem 3.5).
Fix $\epsilon=1/2$. By Proposition B.20, for all $j,N\in\mathbb{N}$, $p=\lceil\log(d)\rceil$, $K\leq\sqrt{d/(8c_{0})}$, we have
@@
with $(P_{i})_{i=1}^{j}$ as in Proposition B.20, $\mu=\mathbb{E}[P_{1}]=(1-\epsilon)K=K/2$. By Hoeffding’s inequality, we have
@@
$=\exp\left(-\frac{2(jK/2-N)^{2}}{jK^{2}}\right)\leq\exp\left(-j\left(1-\frac{2N}{jK}\right)\right)$
for any $j>2N/K$. In particular, for any $\delta>0$, if $j\geq\max\{\frac{4N}{K},2\log(1/\delta)\}$, then $j\left(1-\frac{2N}{jK}\right)\geq 2\log(1/\delta)\left(1-\frac{1}{2}\right)=\log(1/\delta)$ and the above exponential bound is less than $\delta$.
B.6 Proof of Corollary 3.6
Let $(X_{n})_{n=0,1,2,\dots}$ be the RWM trajectory, i.e. the exact solution of the recursion in (1). Under Assumptions 3.1-3.2, $mI_{d}\preccurlyeq\nabla^{2}V$, $h=1$ and $X_{0}\sim\mathcal{N}(x^{\star},L^{-1}I_{d})$, we can apply Theorem 49 and Remark 50 of Andrieu et al. (2024) to deduce $\sqrt{\chi^{2}(\mathcal{L}(X_{n_{mix}}),\pi)}\leq\epsilon$ with $n_{mix}=\mathcal{O}(\frac{L}{m}d)$, where $n_{mix}$ depends on $\epsilon$, $d$, $L/m$ and is explicitly provided in Andrieu et al. (2024, Rmk.50). Consider running the Online Picard algorithm with $K=\sqrt{d/(8c_{0})}$ for
@@
parallel iterations and returning $Y=X^{(J)}_{\min\{n_{mix},L^{(J)}\}}$ as output. When $n_{mix}\leq L^{(J)}$, we have $Y=X^{(J)}_{n_{mix}}=X_{n_{mix}}$. By Theorem 3.5, $\mathbb{P}(Y=X_{n_{mix}})\geq\mathbb{P}(L^{(J)}\geq n_{mix})\geq 1-\epsilon/2$. Therefore
@@
Lemma B.27.
Under Assumption 3.1 with $mI_{d}\preccurlyeq\nabla^{2}V$, we have
$\mathbb{E}[|A_{n}-B_{n}|]$ $\leq 2L(C+d),$ $\hbox{for all }n=1,2,\dots\,$, (17)
$\limsup_{n}\mathbb{P}(|A_{n}|\leq c)$ $=0$ $\hbox{for all }n=1,2,\dots\,,\,c\in(0,\infty).$ (18)
Proof B.28.
Let A=V(x+z)V(x)A=V(x+z)-V(x), B=V(y+z)V(y)B=V(y+z)-V(y), for some $x,y,z\in\mathcal{X}$. By $L$-smoothness of $V$, we have that, for some $t_{1},t_{2}\in(0,1)$ depending on $x,y,z$,
@@
By rotational invariance $Z$, the right hand-side of (19) depends on $\nabla V(x_{n})$ only through its norm. By $m$-convexity of $V$, $\|\nabla V(x_{n})\|\to\infty$ as $|x_{n}|\to\infty$ and, consequently, the right hand-side of (19) goes to 0, proving the desired result in (18).
Lemma B.29.
Under Assumption 3.1 with $mI_{d}\preccurlyeq\nabla^{2}V$, we have
@@
By Lemma B.27, $\limsup_{n\to\infty}\mathbb{P}(A_{n}>c\cap B_{n}>c)=0$ and
@@
which goes to 0 as $n\to\infty$ by Lemma B.29.
B.8 Proof of Proposition 5.1
By Markov inequality and the bound in Theorem 3.1, we have for $K\leq\lfloor bd\rfloor$, $b=\epsilon r/(3c_{0})$,